Show mathematically that the MC curve intersects the AVC curve and the AC curve at the minimum point.

MC=2/3y

AVC=y/3

AC=y/3+6/y
Answer #1

To find the Minimum point of AVC and AC, differentiate them and equate to zero.

dAVC/dy= 1/3=0

So AVC is minimum at y=0

AVC at y=0 is 0

dAC /dy= 1/3-6/y^2=0

1/3= 6/y^2

y^2= 18

y=√18= 4.243

AC is minimum at y= 4.243

AC at y=4.243 is 4.243/3 +6/4.243= 2.828

Now check the value of MC at y=0 and y= 4.243

MC=0/3= 0 = AVC when y=0

MC= 2*4.243/3= 2.828 =AC when y=4.243

So MC intersect AVC and AC at the minimum point